OpenStudy (anonymous):

what is the equation of the line in standard form that passes through the point (-2, -3) and is parallel to the line y= x + 9

OpenStudy (amistre64):

what are some properties the equations of parallel lines ?

OpenStudy (anonymous):

the slope of y=x+9 is 1 use general form y-y1 = m(x-x1) m = slope = 1 , x1=-2, y1=-3 y+3 = x+ 2 rearrange to get your desired equation

OpenStudy (anonymous):

a) x - y = 1 b) x + y = 1 c) x - y = -5 d) x + y = 5

OpenStudy (anonymous):

a